OpenStudy (anonymous):

ile to 2+2*2-2*2-2-2-2*2+2

OpenStudy (anonymous):

14?

OpenStudy (karategirl2002):

14

OpenStudy (mathstudent55):

Start by doing the multiplications.

OpenStudy (anonymous):

-4 is answer XD

OpenStudy (karategirl2002):

14 aculey

OpenStudy (mathstudent55):

2+2*2-2*2-2-2-2*2+2 Start with the multiplications from left to right. = 2 + 4 - 4 - 2 - 2 - 4 + 2 Now do additions and subtractions in the order they appear from left to right. = 6 - 4 - 2 - 2 - 4 + 2 = 2 - 2 - 2 - 4 + 2 = 0 - 2 - 4 + 2 = -2 - 4 + 2 = -6 + 2 = -4

OpenStudy (karategirl2002):

14

OpenStudy (mathstudent55):

@karategirl2002 Since you insist the answer is 14, can you please do it one step at a time, like I did above, to show how you get 14?